# Binary mergers in the centers of galaxies: synergy between stellar flybys and tidal fields

### v1

@@ Abstract
Galactic centers are very dense and dynamically active environments, often harboring a nuclear star cluster and supermassive black hole at their cores. Binaries in these environments are subject to strong tidal fields that can efficiently torque its orbit, exciting near unity eccentricities that ultimately lead to their merger. In turn, the frequent close interactions due to passing stars impulsively perturb the orbit of the binary, generally softening their orbit until their evaporation, thus potentially hindering the role of the tidal fields to drive these mergers. In this work, we study the evolution of compact object binaries in the galactic center and their merger rates, focusing for the first time on the combined effect of the cluster’s tidal field and flyby interactions. We find a significant synergy between both evolution processes, where merger rates increase by a factor of $\sim 10-30$ compared with models in which only flybys or tidal fields are taken into account. This synergy is mainly a consequence of the persistent tides-driven eccentricity excitation that is enhanced by the gradual diffusion of $j_{z}$ ($z$-component of the angular momentum) driven by flybys. The merger efficiency peaks when the diffusion rate is $\sim 10-100$ times slower than the torquing due to the tidal field. Added to this synergy, we also find that the gradual softening of the binary can lift the relativistic quenching of initially tight binaries, otherwise unable to reach extreme eccentricities, and thus expanding the available phase-space for mergers. Cumulatively, we conclude that despite the gradual softening of binaries due to stellar flybys, these greatly enhance their merger rates in the centers of galaxies by promoting the eccentricity excitation driven by the tidal fields.
@@ 1 Introduction
Most nearby galaxies are known to harbour very dense stellar clusters at their core (Neumayer et al., 2011; Turner et al., 2012; Georgiev & Böker, 2014). These nuclear star clusters (NSCs) tend to have masses of roughly $10^{5}-10^{8}M_{\odot}$, a small effective radii of only a few parsecs, and usually have a supermassive black hole (SMBH) residing at their center. This is the case for our own galaxy, containing an NSC with a mass of $\sim 3\times 10^{7}M_{\odot}$ (Schödel et al., 2014, 2018) and the central SMBH $*{}^{*}$ with a mass of about $4\times 10^{6}M_{\odot}$ (Ghez et al., 2005; Gillessen et al., 2009). This makes the galactic center (GC) a very dense and dynamical environment, which turns it into an ideal laboratory to test several astrophysical phenomena involving stellar dynamics, general relativity (GR), and gravitational waves (GWs).
In the last decade there has been an increasing interest in gravitational wave astrophysics, due to GW detections from the LIGO and VIRGO interferometers. Compact object binary mergers are known to be one of the main GW emitting events (Abbott et al., 2016, 2017), therefore recent studies have been focused on the dynamical mechanisms that lead to these mergers in order to better understand GW sources. Particularly in GCs, one of the most important physical processes leading up to mergers is the torquing of binary’s orbits produced by the SMBH, process known as the von Zeipel-Lidov-Kozai (ZKL) mechanism (Antonini & Perets, 2012; Stephan et al., 2019; Hoang et al., 2019; Petrovich & Antonini, 2017). More generally, we can consider the tidal field from both the SMBH and the cluster which produces a behaviour similar to the ZKL mechanism (Petrovich & Antonini, 2017; Hamilton & Rafikov, 2019; Bub & Petrovich, 2020; Hamilton & Rafikov, 2021).
@@
Due to the high velocity dispersion $\sigma$ in the GC, most binaries in the GC are soft (e.g., $\sigma$ is larger than their Keplerian velocity). One single encounter can potentially reset a binary’s orbit, changing its orientation and even pushing it towards extreme eccentricities. The semi-major axis of these binaries will also tend to increase due to these encounters (Heggie, 1975; Hills, 1975a, b). This orbit widening limits the efficiency for tidal field driven mergers, and previous works have simply prescribed it as a limiting evaporation timescale $t_{\text{evap}}$ (e.g., Stephan et al. 2016; Petrovich & Antonini 2017; Hoang et al. 2018), using $t_{\text{evap}}$ from Binney & Tremaine (2008). However, beyond the orbit softening, flyby interactions can produce a random walk in the binary’s angular momentum $\vec{j}$ which may lead to mergers as studied by Michaely & Perets (2020). A study looking into the effect from flybys in the cluster-driven dynamics is largely missing in this context (GCs) and the focus of our work.
@@ 2 Dynamical evolution due to tidal fields
As binaries are usually immersed in stellar clusters, an important aspect of their dynamical evolution will be driven by the smooth tidal field of said cluster. These tidal fields drive secular oscillations in a binary’s eccentricity and inclination, similar to the ZKL mechanism. In this section we model this evolution mathematically, following the work done by Hamilton & Rafikov (2019a, b).
@@ 2.1 Equations of motion
As shown by Hamilton & Rafikov (2019a), in this torus-averaged limit the tidal tensor of a given potential $\Phi$ only has diagonal terms, and $\langle\Phi_{xx}\rangle=\langle\Phi_{yy}\rangle$. For a smooth potential $\Phi$ that changes over scales that are much larger than that of the inner binary’s semi-major axis (expanded up to the quadrupole tidal field), this results in the torus-averaged potential
@@
$\begin{split}H_{1}=2\Bigl{[}1-\frac{3}{2}e^{2}(5\Gamma-1)-3\Gamma(\vec{j}\cdot\hat{j}_{\text{out}})^{2}+15\Gamma(\vec{e}\cdot\hat{j}_{\text{out}})^{2}\Bigr{]},\end{split}$ (3)
where $\{\vec{e},\vec{j}\}$ are Milankovitch’s eccentricity and dimensionless angular momentum vectors (e.g., see Tremaine 2023). The associated dimensionless time is $\tau=t/\tau_{\text{sec}}$, where $\tau_{\text{sec}}$ is the secular timescale given by
@@
As discussed by Hamilton & Rafikov (2019b, a), $\Gamma$ determines the phase-space structure of the Hamiltonian, and hence is a useful parameter to classify different dynamical regimes for binary evolution. For example, $\Gamma=1$ when the potential is Keplerian (only a SMBH at the center), and hence the Hamiltonian in equation (3) reduces to the quadrupole ZKL Hamiltonian. The secular timescale reduces to the well-known ZKL timescale (e.g., Holman et al. 1997):
$\tau_{\rm ZKL}$ $=$ $\frac{4a^{3}_{\text{out}}(1-e_{\text{out}}^{2})^{3/2}M^{1/2}_{\text{bin}}}{3a_{\text{in}}^{3/2}M_{\text{BH}}G^{1/2}}$
@@ 2.2 Merger conditions
$\tau_{\text{GW}}=\frac{3}{85}\left(\frac{a_{\text{in}}^{4}c^{5}}{G^{3}m_{1}m_{2}M_{\text{bin}}}\right)(1-e_{\text{in}})^{7/2}.$ (12)
If a binary’s orbit shrinks significantly in one secular cycle, then a rapid merger will take place. For this to happen, the timescale for the change in the semi-major axis of the binary due to GW losses ($\tau_{\text{GW}}$) must be shorter than the timescale for the change in the pericentric distance $a_{\text{in}}(1-e_{\text{in}})$ due to tidal fields, which, at high eccentricities, is $\simeq\sqrt{1-e^{2}}\tau_{\rm sec}$. This results in the condition
@@
for binaries with $M_{\text{bin}}=10M_{\odot}$ initially at $a_{\text{in}}=10$au when only ZKL cycles are considered (only SMBH). A similar expression can be obtained for a more general cluster potential (arbitrary $\Gamma$; see Equation 16 for a secular timescale adding a cluster with a Hernquist potential).
@@ 3 Dynamics in the Galactic center
For small values of $a_{\text{out}}/s$ we have $\Gamma\approx 1$, where the SMBH dominates over the cluster and hence the dynamical evolution of the binary reduces to the ZKL mechanism as mentioned in section §2.1. The same happens when $a_{\text{out}}/s$ becomes very large; the cluster and SMBH in this limit behave as a single body so that as we get further away from the central SMBH, $\Gamma\rightarrow 1$.
@@
Figure 1: Full evolution of a binary due to both tidal fields and flyby effects. The initial parameters are $a_{\text{out}}=0.3$ pc, $a_{\text{in}}=25$ AU, $e_{\text{in}}=0.1$, $\omega_{\text{in}}=\pi/4$, $\Omega_{\text{in}}=0$, $i_{\text{in}}=80$°. Each stellar component of the binary has a mass of $5M_{\odot}$, and we consider the fiducial scenario of $M_{\text{BH}}=4\times 10^{6}M_{\odot}$. The red curves represent the evolution in a flyby free scenario, by simply integrating Equations (5) and (6). In panels (d) and (e), the shaded region contains the median and both upper and lower quartiles.
@@ 4 Numerical simulations
[…]
@@ 4.1 Population statistics
In Figure 2 we show the CDF of the maximum eccentricities achieved by a population of $\sim 700$ binaries in the galactic center at different stages of their evolution. The initial eccentricities of these binaries follow a thermal distribution such that $f(e)de=2ede$, and their initial inclinations are uniform such that $\cos{i_{0}}$ is also uniform between -1 and 1. The eccentricity of the outer orbits is set to 0. All binaries are twins with a total mass of $10M_{\odot}$ and are evolved for $100\tau_{0}$, $\tau_{0}$ being the initial secular timescale of each individual binary according to equation (16). We consider the fiducial scenario of our GC, where $M_{\text{BH}}=4\times 10^{6}M_{\odot}$, $M_{\text{cl}}\approx 8M_{\text{BH}}$, and $s=4$ pc.
@@
In Figure 3 we can see the CDFs of the maximum eccentricities achieved after $100\tau_{0}$ by three different populations of binaries. One population was evolved considering only tidal field effects, the second considering only flybys and the third taking into account both effects (the same population as shown in figure 2). The figure shows that tidal fields are the least efficient dynamical mechanism in producing extreme eccentricities; only about $\sim 1\%$ of binaries evolved with tidal fields alone reach $1-e_{\text{max}}<10^{-4}$. Flybys on the other hand prove to be slightly more effective, driving about $3\%$ of binaries in the population to said values. However, when combining both mechanisms a clear synergy is found, increasing this probability up to $\sim 34\%$. For reference, the bottom panel shows the minimum pericenter distances for the same population of binaries. Although several binaries reach very high eccentricities, as was mentioned in section §3.1 these binaries will tend to get softer, which means that $a_{\text{in}}$ and hence the pericentric distance $q=a_{\text{in}}(1-e_{\text{in}})$ will grow as well. This is where the stability criterion mentioned previously in equation (20) becomes important. As can be seen in the CDF, the extreme values of the eccentricities achieved seem to dominate over the increase in $a_{\text{in}}$, producing a considerable amount of binaries with pericentric distances smaller than $\sim 10^{-3}$au, equivalent to a merger timescale $\tau_{\rm GW}<3\times 10^{5}$ yrs (Eq. [12]). The latter may be shorter than the ZKL timescale in Equation (2.1) and, therefore, expected to lead to a merger within one secular cycle.
@@ 5.1 Experiments changing $\mathcal{R}$
In Figure 6 we show the merger fractions of different binary populations as a function of $\mathcal{R}$, which was varied by changing the value of $M_{\text{BH}}$. These populations were all integrated for $\sim 15$ Myr, which is roughly $50\tau_{\text{sec}}$ for a binary starting with $a_{\text{in}}=25$au at $0.3$pc considering the fiducial case of $\mathcal{R}=0.02$. We consider any binary that reaches $1-e_{\text{max}}\leq 10^{-4}$ to be a merger, as was discussed in section §2.2. As mentioned previously, a less diffusive regime comes with shorter secular timescales, therefore by integrating a population with $\mathcal{R}=0.006$ for $15$Myr we are actually letting them evolve for $\sim 500\tau_{\text{sec}}$. However as we keep lowering $\mathcal{R}$ by increasing $M_{\text{BH}}$, binaries start to become unstable very fast as according to equation (20) and hence the likelihood of them merging before they evaporate decreases. We also move into a regime in which the binaries only experience a few flyby interactions in each cycle, which could also contribute in decreasing the probability of reaching extreme eccentricity values as occurs in figure 6. Our experiments suggest that there is an optimal value for $\mathcal{R}\sim 0.02$ at which the merger fraction peaks.

### v2

@@ Abstract
Galactic centers are very dynamically active environments, often harbouring a nuclear star cluster and supermassive black hole at their cores. Binaries in these environments are subject to strong tidal fields that can efficiently torque its orbit, exciting near unity eccentricities that ultimately lead to their merger. In turn, frequent close interactions with passing stars impulsively perturb the orbit of the binary, generally softening their orbit until their evaporation, potentially hindering the role of tides to drive these mergers. In this work, we study the evolution of compact object binaries in the galactic center and their merger rates, focusing for the first time on the combined effect of the cluster’s tidal field and flyby interactions. We find a significant synergy between both processes, where merger rates increase by a factor of $\sim 10-30$ compared to models in which only flybys or tides are taken into account. This synergy is a consequence of the persistent tides-driven eccentricity excitation that is enhanced by the gradual diffusion of $j_{z}$ driven by flybys. The merger efficiency peaks when the diffusion rate is $\sim 10-100$ slower than the tides-driven torquing. Added to this synergy, we also find that the gradual softening of the binary can lift the relativistic quenching of initially tight binaries, otherwise unable to reach extreme eccentricities, and thus expanding the available phase-space for mergers. Cumulatively, we conclude that despite the gradual softening of binaries due to flybys, these greatly enhance their merger rates in galaxy centers by promoting the tidal field-driven eccentricity excitation.
@@ 1 Introduction
Most nearby galaxies are known to harbor very dense stellar clusters at their core (Neumayer et al., 2011; Turner et al., 2012; Georgiev & Böker, 2014). These nuclear star clusters (NSCs) tend to have masses of roughly $10^{5}-10^{8}M_{\odot}$, a small effective radii of only a few parsecs, and usually have a supermassive black hole (SMBH) residing at their center. This is the case for our own galaxy, containing an NSC with a mass of $\sim 3\times 10^{7}M_{\odot}$ (Schödel et al., 2014, 2018) and the central SMBH SgrA with a mass of about $4\times 10^{6}M_{\odot}$ (Ghez et al., 2005; Gillessen et al., 2009). This makes the galactic center (GC) a very dense and dynamical environment, which makes it an ideal laboratory to test several astrophysical phenomena involving stellar dynamics, general relativity (GR), and gravitational waves (GWs).
In the last decade there has been an increasing interest in gravitational-wave astrophysics, due to GW detections from the LIGO and VIRGO interferometers. Compact object binary mergers are known to be one of the main GW emitting events (Abbott et al., 2016, 2017), therefore recent studies have focused on the dynamical mechanisms that lead to these mergers to better understand the sources of GWs. Particularly in GCs, one of the most important physical processes leading up to mergers is the torquing of binary’s orbits produced by the SMBH, process known as the von Zeipel-Lidov-Kozai (ZLK) mechanism (Antonini & Perets, 2012; Stephan et al., 2019; Hoang et al., 2019; Petrovich & Antonini, 2017). More generally, we can consider the tidal field from both the SMBH and the cluster which produces a behaviour similar to the ZLK mechanism (Petrovich & Antonini, 2017; Hamilton & Rafikov, 2019; Bub & Petrovich, 2020; Hamilton & Rafikov, 2021).
@@
Due to the high velocity dispersion $\sigma$ in the GC, most binaries in the GC are soft (e.g., $\sigma$ is larger than their Keplerian velocity) (Stephan et al., 2016; Alexander & Pfuhl, 2014; Rose et al., 2020; Gautam et al., 2024). One single encounter can potentially reset a binary’s orbit, changing its orientation and even pushing it towards extreme eccentricities. The semi-major axis of these binaries will also tend to increase due to these encounters (Heggie, 1975; Hills, 1975a, b). This orbit widening limits the efficiency for tidal field driven mergers, and previous works have simply prescribed it as a limiting evaporation timescale $t_{\text{evap}}$ (e.g., Stephan et al. 2016; Petrovich & Antonini 2017; Hoang et al. 2018), using $t_{\text{evap}}$ from Binney & Tremaine (2008). However, beyond the orbit softening, flyby interactions can produce a random walk in the binary’s angular momentum $\vec{j}$ which may lead to mergers as studied by Michaely & Perets (2020). A study looking into the effect from flybys in the cluster-driven dynamics is largely missing in this context (GCs) and the focus of our work.
@@ 2 Dynamical evolution due to tidal fields
As binaries are usually immersed in stellar clusters, an important aspect of their dynamical evolution will be driven by the smooth tidal field of said cluster. These tidal fields drive secular oscillations in a binary’s eccentricity and inclination, similar to the ZLK mechanism. In this section we model this evolution mathematically, following the work done by Hamilton & Rafikov (2019a, b). We will refer to the compact object binary system as the inner binary, while the orbit of this system around the central SMBH will be referred to as the outer binary.
@@ 2.1 Equations of motion
As shown by Hamilton & Rafikov (2019a), in this torus-averaged limit the tidal tensor of a given potential $\Phi$ only has diagonal terms, and $\langle\Phi_{xx}\rangle=\langle\Phi_{yy}\rangle$. Here it is useful to introduce Milankovitch’s eccentricity and dimensionless angular momentum vectors $\{\vec{e},\vec{j}\}$, Milankovitch’s eccentricity and dimensionless angular momentum vectors (e.g., see Tremaine 2023), in order to simplify notation. For a smooth potential $\Phi$ that changes over scales that are much larger than that of the inner binary’s semi-major axis (expanded up to the quadrupole tidal field), this results in the torus-averaged potential
@@
$\begin{split}H_{1}=2\Bigl{[}1-\frac{3}{2}e^{2}(5\Gamma-1)-3\Gamma(\vec{j}\cdot\hat{j}_{\text{out}})^{2}+15\Gamma(\vec{e}\cdot\hat{j}_{\text{out}})^{2}\Bigr{]}.\end{split}$ (3)
The associated dimensionless time is $\tau=t/\tau_{\text{sec}}$, where $\tau_{\text{sec}}$ is the secular timescale given by
@@
As discussed by Hamilton & Rafikov (2019b, a), $\Gamma$ determines the phase-space structure of the Hamiltonian, and hence is a useful parameter to classify different dynamical regimes for binary evolution. For example, $\Gamma=1$ when the potential is Keplerian (only a SMBH at the center), and hence the Hamiltonian in equation (3) reduces to the quadrupole ZLK Hamiltonian. The secular timescale reduces to the well-known ZLK timescale (e.g., Holman et al. 1997):
$\tau_{\rm ZLK}$ $=$ $\frac{4a^{3}_{\text{out}}(1-e_{\text{out}}^{2})^{3/2}M^{1/2}_{\text{bin}}}{3a_{\text{in}}^{3/2}M_{\text{BH}}G^{1/2}}$ (7)
@@
We note that the ZLK Hamiltonian assumes the test particle approximation and sets $e_{\text{out}}=0$, so that the quadrupole level expansion is a good approximation of the secular dynamics. If, instead, the outer orbit is eccentric, the $z$-component of the angular momentum is no longer conserved, and hence the octupole expansion is required (Naoz, 2016). This expansion allows for qualitatively different behavior, where extremely high eccentricities for the inner orbit can be achieved along with orbital flipping (Naoz et al., 2011, 2013; Hoang et al., 2018).
@@ 2.2 Merger conditions
$\tau_{\text{GW}}=\frac{3}{85}\left(\frac{a_{\text{in}}^{4}c^{5}}{G^{3}m_{1}m_{2}M_{\text{bin}}}\right)(1-e_{\text{in}})^{7/2},$ (11)
where $m_{1}$ and $m_{2}$ If a binary’s orbit shrinks significantly in one secular cycle, then a rapid merger will take place. For this to happen, the timescale for the change in the semi-major axis of the binary due to GW losses ($\tau_{\text{GW}}$) must be shorter than the timescale for the change in the pericentric distance $a_{\text{in}}(1-e_{\text{in}})$ due to tidal fields, which, at high eccentricities, is $\simeq\sqrt{1-e^{2}}\tau_{\rm sec}$. This results in the condition
@@
for binaries with $M_{\text{bin}}=10M_{\odot}$ initially at $a_{\text{in}}=10$au when only ZLK cycles are considered (only SMBH). A similar expression can be obtained for a more general cluster potential (arbitrary $\Gamma$; see Equation 15 for a secular timescale adding a cluster with a Hernquist potential).
@@ 3 Dynamics in the Galactic center
For small values of $a_{\text{out}}/s$ we have $\Gamma\approx 1$, where the SMBH dominates over the cluster and hence the dynamical evolution of the binary reduces to the ZLK mechanism as mentioned in section §2.1. The same happens when $a_{\text{out}}/s$ becomes very large; the cluster and SMBH in this limit behave as a single body so that as we get further away from the central SMBH, $\Gamma\rightarrow 1$.
@@
Figure 1: Full evolution of a binary due to both tidal fields and flyby effects. The initial parameters are $a_{\text{out}}=0.3$ pc, $a_{\text{in}}=25$ AU, $e_{\text{in}}=0.1$, $\omega_{\text{in}}=\pi/4$, $\Omega_{\text{in}}=0$, $i_{\text{in}}=80$°. Each stellar component of the binary has a mass of $5M_{\odot}$, and we consider the fiducial scenario of $M_{\text{BH}}=4\times 10^{6}M_{\odot}$. The red curves represent the evolution in a flyby-free scenario, simply by integrating Equations (5) and (6). In panels (d) and (e), the shaded region contains the median and both upper and lower quartiles.
@@ 3.1 Fly-bys
The hard-soft boundary then is given by $v_{\rm{kep}}=\sigma$. For example, if we consider a twin binary with two $5M_{\odot}$ components at $a_{\text{out}}=0.3$pc , this boundary would be given by $a_{\text{in}}=0.2$au. If we increase the mass of the binary component this limit will occur at a larger value; for $10M_{\odot}$ components it occurs at $a_{\text{in}}=0.4$au, and for $50M_{\odot}$ at $a_{\text{in}}=2.1$au. These continue to be relatively small values, further reinforcing that most binaries in the galactic center can be classified as soft.
@@ 4 Numerical simulations
We establish the velocity of our flybys to be constant at $v_{p}=200$ kms-1, considering that within 1 pc of the center of our galaxy $\sigma$ varies roughly between 100 and 300 kms-1 as according to equation (16). These simplifications allow for controlled simulations, which makes comparisons with analytical estimates such as the ones made in section §5 easier.
@@ 4.1 Population statistics
In Figure 2 we show the CDF of the maximum eccentricities achieved by a population of $\sim 700$ binaries in the galactic center at different stages of their evolution. The initial eccentricities of these binaries follow a thermal distribution such that $f(e)de=2ede$, and their initial inclinations are uniform such that $\cos{i_{0}}$ is also uniform between -1 and 1. The eccentricity of the outer orbits is set to 0. All binaries are twins with a total mass of $10M_{\odot}$ and are evolved for $100\tau_{0}$, $\tau_{0}$ being the initial secular timescale of each individual binary according to equation (15). We consider the fiducial scenario of our GC, where $M_{\text{BH}}=4\times 10^{6}M_{\odot}$, $M_{\text{cl}}\approx 8M_{\text{BH}}$, and $s=4$ pc. The initial values for both $a_{\text{in}}$ and $a_{\text{out}}$ are drawn from a uniform distribution, ranging between $15-50$au for $a_{\text{in}}$ and $0.1-1$pc for $a_{\text{out}}$. This means that the total population of binaries is initially soft, and that the velocity dispersion does not exceed $275$kms-1 making our assumption of $v_{p}=200$kms-1 appropriate.
@@
In Figure 3 we can see the CDFs of the maximum eccentricities achieved after $100\tau_{0}$ by three different populations of binaries. One population was evolved considering only tidal field effects, the second considering only flybys and the third taking into account both effects (the same population as shown in figure 2). The figure shows that tidal fields are the least efficient dynamical mechanism in producing extreme eccentricities; only about $\sim 1\%$ of binaries evolved with tidal fields alone reach $1-e_{\text{max}}<10^{-4}$. Flybys on the other hand prove to be slightly more effective, driving about $3\%$ of binaries in the population to said values. However, when combining both mechanisms a clear synergy is found, increasing this probability up to $\sim 34\%$. For reference, the bottom panel shows the minimum pericenter distances for the same population of binaries. Although several binaries reach very high eccentricities, as was mentioned in section §3.1 these binaries will tend to get softer, which means that $a_{\text{in}}$ and hence the pericentric distance $q=a_{\text{in}}(1-e_{\text{in}})$ will grow as well. This is where the stability criterion mentioned previously in equation (19) becomes important. As can be seen in the CDF, the extreme values of the eccentricities achieved seem to dominate over the increase in $a_{\text{in}}$, producing a considerable amount of binaries with pericentric distances smaller than $\sim 10^{-3}$au, equivalent to a merger timescale $\tau_{\rm GW}<3\times 10^{5}$ yrs (Eq. [11]). The latter may be shorter than the ZLK timescale in Equation (7) and, therefore, expected to lead to a merger within one secular cycle.
@@ 5.1 Experiments changing $\mathcal{R}$
In Figure 6 we show the merger fractions of different binary populations as a function of $\mathcal{R}$, which was varied by changing the value of $M_{\text{BH}}$. These populations were all integrated for $\sim 15$ Myr, which is roughly $50\tau_{\text{sec}}$ for a binary starting with $a_{\text{in}}=25$au at $0.3$pc considering the fiducial case of $\mathcal{R}=0.02$. We consider any binary that reaches $1-e_{\text{max}}\leq 10^{-4}$ to be a merger, as was discussed in section §2.2. As mentioned previously, a less diffusive regime comes with shorter secular timescales; therefore, by integrating a population with $\mathcal{R}=0.006$ for $15$Myr we are actually letting them evolve for $\sim 500\tau_{\text{sec}}$. However, as we keep lowering $\mathcal{R}$ by increasing $M_{\text{BH}}$, binaries start to become unstable very fast according to equation (19) and hence the likelihood of them merging before they evaporate decreases. We also move into a regime in which the binaries only experience a few flyby interactions in each cycle, which could also contribute to decreasing the probability of reaching extreme eccentricity values as occurs in figure 6. Our experiments suggest that there is an optimal value for $\mathcal{R}\sim 0.02$ at which the merger fraction peaks.
5.2 Understanding the optimal value of $\mathcal{R}$ to increase the merger fractions
We provide an analytical explanation at the order-of-magnitude level of why there is a particular value for $\mathcal{R}$ at which the merger fraction peaks as shown in figure 6.
We can write our merger condition as $j_{\text{min}}=\sqrt{1-e_{\text{max}}^{2}}$ reaching values below a critical specific angular momentum $j_{\text{crit}}$. The latter is set to $j_{\text{crit}}\simeq 0.02$ in our simulations ($1-e_{\rm crit}=10^{-4}$). In the absence of flybys ($\mathcal{R}=0$), the fraction of mergers for initially circular and isotropic orbits is simply [^1] The merger fraction is given by the probability that $j_{\text{min}}=\sqrt{5/3}|\cos I|<j_{\rm crit}$, which becomes $\sqrt{3/5}j_{\rm crit}$ for an isotropic distribution. $\sqrt{3/5}j_{\rm crit}\simeq 0.015$ in our simulations. This changes slightly to $\simeq 0.52j_{\rm crit}\simeq 0.01$ for an initial thermal distribution.
As Figure 4 shows, the dominant effect of flybys on the merger rates comes from changes in inclinations, which corresponds to changes in $j_{z}$ at fixed $j$. Binaries that reach $|j_{z}|<j_{\text{crit}}$ can merge, provided that the binary lies in that range for at least a significant fraction of a secular timescale so it can be assisted by the tidal field. In a completely diffusive scenario, $j_{z}$ undergoes a random walk due to stellar encounters such that $\langle j_{z}^{2}\rangle=\mathcal{D}t$, with $\mathcal{D}$ a constant diffusive parameter. Thus, in a relaxation timescale $1/\mathcal{D}$ binaries outside the range $|j_{z}|<j_{\text{crit}}$ will be kicked in, monotonically increasing the fraction of mergers with $\mathcal{R}\propto\mathcal{D}$ consistent with the increase in Figure 6 below $\mathcal{R}\sim 0.02$.
As we continue to increase the diffusion rate due to stellar encounters $\mathcal{D}$, the merger fraction is expected to turn over when the binaries are kicked in and out of the $|j_{z}|<j_{\text{crit}}$ range in timescale shorter than the secular time. The critical diffusion timescale $t_{\text{crit}}$ to linger in this region is given by
$t_{\text{crit}}\sim j_{\text{crit}}^{2}/\mathcal{D}.$ (23)
Thus, we expect flybys to most efficiently boost the merger rate when $t_{\text{crit}}\sim\tau_{0}$, a condition that can be expressed in terms of our definition of $\mathcal{R}$ as
$\mathcal{R}=(\tau_{0}\mathcal{D})^{1/2}\sim j_{\text{crit}},$ (24)
which would be the value of $\mathcal{R}$ for which the merger fraction peaks. In our simulations, this estimate predicts a peak value at $\mathcal{R}\sim 0.02$, which perfectly matches the observed peak in Figure 6.
However, it is important to note that our simulations are not in a fully diffusive regime and we observe sudden large jumps in $a_{\text{in}}$ (see Figure 5) due to penetrative encounters ($b<a_{\text{in}}$). Therefore, Equation (24) serves as a general approximation to where the merger fraction should peak, but can be slightly off due to the non-diffusive nature of these penetrative flybys. In order to illustrate the approximate nature of our estimates, we repeat the numerical experiments for different values of $j_{\rm crit}$ in Figure B1. We observe that even though there is a value for $\mathcal{R}$ at which the merger fraction peaks, and it decreases for lower values of $j_{\text{crit}}$, the peaks do not occur exactly at $\mathcal{R}\sim j_{\text{crit}}$.